HUBBLE SPACE TELESCOPE DAILY REPORT #1921 PERIOD COVERED: 0000Z (UTC) 07/16/97 - 0000Z (UTC) 07/17/97 Daily Status Report as of 198/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of dark calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled with no problems. 1.4 Completed Two Sets of STIS/CCD/MA2 7087 (OTA-STIS Pointing and Throughput Stability) The Space Telescope Imaging Spectrograph (Both CCD and MA2) were used to make observations of stars HD192273 and HD177989 in order to calibrate the OTA-STIS pointing and throughput stability. This is to evaluate the image motion at both slit and detector planes in a thermally stable case and through a relatively severe thermal slew on both second (jitter) and hour (drift) time scales. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrently with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Operations Notes: The NSSC-1 load uplinked first at 197/184907Z experienced several problems. On the first attempt an EXEC 101 error and a count failure occurred. The re-transmission saw a count failure and a CU/SDF parity error. After the parity error was reset by ground command, the load was successfully uplinked. HSTAR 6124 was written. The SSR Interrupt Log was dumped at 197/2242Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH